OpenStudy (anonymous):

the median size of a new single family home grew from 2174 ft^2 in 1998 to 2607 ft^2 in 2009. what is the percetn increase?

OpenStudy (ladiesman217):

i got 2%\[2607-2174=x\] \[2009-1998=y\] what are x and y

OpenStudy (anonymous):

the percent of increase is about what percent?

OpenStudy (ladiesman217):

im showing you how to get it 2009−1998=11 2607−2174=433 11/433=2.5

OpenStudy (anonymous):

thanks ladiesman217

OpenStudy (ladiesman217):

no problem